OpenStudy (anonymous):

Evalute 2x3 - 3x2 + 4x + 5 with x = 4.

OpenStudy (anonymous):

2x3 - 3x2 + 4x + 5 with x = 4. 2x^3 - 3x^2 +4x + 5 = 2 (4^3) - 3(4^2) + 4(4) = (2xx64) - (3xx16) + 14 = 128 - 48 + 14 = 128 - 48 + 14 + 5 = 101